# AutoV: Learning to Retrieve Visual Prompt for Large Vision-Language Models

### v1

@@ 4.2 Main Results
| Method | MMMU | POPE | VizWiz | MMVet | VQA | LLaVA | Avg. |
| LLaVA-v1.5 (7B) | |||||||
| Base Model | 36.3 | 85.9 | 50.0 | 30.6 | 58.2 | 65.4 | |
| FGVP (NIPS23) | 36.9 | 72.6 | 47.5 | 29.0 | 46.3 | 55.6 | |
| Circle (ICCV23) | 37.0 | 84.5 | 49.6 | 31.3 | 55.5 | 62.2 | |
| API (ECCV24) | 37.4 | 86.0 | 50.3 | 31.8 | 57.7 | 65.8 | |
| AutoV | 38.3 | 86.8 | 51.0 | 32.8 | 58.8 | 66.6 | |
| LLaVA-v1.5 (13B) | |||||||
| Base Model | 36.7 | 85.9 | 53.6 | 36.1 | 61.2 | 66.6 | |
| FGVP (NIPS23) | 36.9 | 66.8 | 46.8 | 29.7 | 49.2 | 63.7 | |
| Circle (ICCV23) | 36.9 | 84.6 | 53.8 | 31.2 | 58.1 | 66.9 | |
| API (ECCV24) | 37.3 | 85.6 | 54.6 | 35.6 | 60.6 | 67.6 | |
| AutoV | 38.5 | 86.9 | 55.4 | 37.0 | 61.9 | 68.5 | |
| LLaVA-OneVision (7B) | |||||||
| Base Model | 49.4 | 89.1 | 58.2 | 48.8 | 76.1 | 80.6 | |
| FGVP (NIPS23) | 48.6 | 82.9 | 52.0 | 45.5 | 65.4 | 70.8 | |
| Circle (ICCV23) | 48.4 | 88.2 | 59.6 | 47.2 | 73.0 | 76.9 | |
| API (ECCV24) | 49.0 | 88.9 | 58.7 | 48.1 | 75.9 | 80.9 | |
| AutoV | 50.0 | 90.0 | 59.4 | 49.9 | 77.3 | 82.3 | |
| InternVL2 (8B) | |||||||
| Base Model | 48.6 | 87.8 | 58.6 | 46.5 | 77.0 | 74.6 | |
| FGVP (NIPS23) | 47.2 | 84.2 | 52.3 | 36.8 | 70.7 | 70.2 | |
| Circle (ICCV23) | 48.3 | 86.0 | 57.5 | 40.3 | 74.9 | 74.0 | |
| API (ECCV24) | 48.3 | 87.9 | 58.0 | 46.3 | 77.0 | 75.6 | |
| AutoV | 49.2 | 88.5 | 59.1 | 47.2 | 77.9 | 76.8 | |
| Qwen2.5-VL (7B) | |||||||
| Base Model | 50.1 | 87.5 | 69.5 | 56.6 | 83.0 | 98.0 | |
| FGVP (NIPS23) | 49.6 | 81.6 | 65.8 | 43.1 | 78.0 | 90.4 | |
| Circle (ICCV23) | 48.2 | 84.2 | 66.7 | 44.5 | 80.5 | 89.5 | |
| API (ECCV24) | 51.0 | 87.6 | 69.3 | 56.8 | 82.3 | 98.1 | |
| AutoV | 52.0 | 88.1 | 70.0 | 57.6 | 83.4 | 98.8 | |
Our results are listed in Table 1, and we compare AutoV with three previous visual prompting methods designed for LVLMs, including FGVP [53], RedCircle [43], and API [55]. Performance improvements and decrements listed in the table are calculated relative to the base model. The training data and implementation details can be found in Appendix B and D.
@@
The main observations from the experimental results are as follows: (1) AutoV significantly outperforms existing methods, demonstrating consistent advantages across all models and benchmarks. Specifically, for LLaVA-7B, LLaVA-13B, LLaVA-OneVision, InternVL2, and Qwen2.5-VL, the average improvements relative to the base model are , , , and , respectively. Our method performs particularly well for LLaVA-7B on MMVet, with an improvement of . Excluding it, our method appears more effective and robust on the VQA task, where other methods all fail to achieve positive gain. (2) Our approach shows the most significant performance advantage on MMVet compared to the API. With an average score of for our method versus for theirs, the relative difference of underscores the importance of accurate case-level visual cues retrieval instead of benchmark-level visual prompts selection. The other two visual prompting methods, lacking the ability to adapt to different queries, do not perform well on VQA tasks. (3) Our method further supplements inherently less accurate models. For instance, LLaVA-7B equipped with our approach, achieves higher accuracy on the MMMU and POPE tasks than the 13B-scale model. This means that our method can be utilized for model efficiency on certain specific tasks: a few additional MLP layers are a better choice compared with hundreds of times larger in parameter size. (4) Both Qwen2.5-VL and InternVL2 integrate the retrieval-augmented prompting strategy from AutoV, which is pre-trained on LLaVA-OneVision. This integration results in an average accuracy gain of . Notably, the improvements are consistent and observed in all individual tasks, including complex reasoning (e.g., MMVet) and real-world scenarios (e.g., VizWiz). This demonstrates that the effectiveness of the retrieval-enhanced prompting strategy is model-agnostic and generalizes well across heterogeneous LVLM architectures, regardless of pretraining scales or backbone variations.
4.3 Ablation Study
4.3.1 How does ranking compare to other retrieval methods?
@@
| only | 36.5 | 50.0 | 31.8 | 39.4 |
| only | 36.9 | 50.3 | 31.2 | 39.5 |
| only | 37.4 | 49.9 | 31.7 | 39.7 |
| only | 37.1 | 50.3 | 30.8 | 39.7 |
| Random | 37.2 | 50.4 | 31.6 | 39.7 |
| Mixture-of-Experts | 37.6 | 50.4 | 31.7 | 39.9 |
| AutoV | 38.3 | 51.0 | 32.8 | 40.7 |
@@
To better analyze why ranking is more suitable for retrieving visual prompts, we list three other retrieval methods and provide experimental evidence for comparison, including heuristic retrieval, random retrieval, and mixture-of-experts retrieval. All experiments are conducted on LLaVA-v1.5.
@@
Heuristic Fixed Retrieval. Heuristic retrieval only involves a single fixed type of visual prompts, e.g., red circle [43], or attention image [55], based on its predefined rules or heuristics. Furthermore, to identify the best rules, they need to conduct benchmark-level multiple rounds of trials, consuming a significant amount of reasoning resources. As shown in rows 3 to 6 of Table 2, we find that each visual prompt could potentially be the optimal solution for different benchmarks. For instance, achieves the best results on MMMU, while is optimal on MMVet, and there is a significant variation in the performance of different types of visual prompts across various benchmarks. In summary, heuristic retrieval offers simplicity in retrieval processes with the need for conducting benchmark-level multiple rounds of trials.
@@
Random Retrieval. Random retrieval picks up one visual prompt from the candidate pool without any specific criteria or order. The th row of Table 2 lists its results, comparing with the individual visual prompt, we observe that the random retrieval’s performance is on par with or slightly superior across the three benchmarks. For instance, the random algorithm outperforms in MMMU by , in MMVet by , and even surpasses the best results of the VP candidates in VizWiz. The above experiment demonstrates that by flexibly selecting from multiple visual prompt candidates, even random retrieval has the potential to outperform individual visual prompts, and the conclusion provides valuable insights for our approach.
@@
MoE (or GateNet) Retrieval. Mixture-of-experts (MoE) retrieval is a prevalent technique used to identify the most relevant expert for a model decision-making process, where the Gumbel Softmax function is employed to assign weights to the GateNet in a differentiable manner during training [16]. In our scenario, the features of multiple visual prompts are input to GateNet to determine which one should be retained and fed into the LLM. We insert GateNet into the output of the projection layer and train it with the same amount of data and training time as AutoV. The results can be seen in the th row of Table 2. We observe a improvement with MoE compared to the baseline. However, the improvement compared to the random algorithm was only , lagging significantly behind AutoV by . We suspect that the Mixture-of-Experts architecture encounters a lack of generalization ability during the fine-tuning stage, which can easily lead to overfitting. Additionally, this architecture does not incorporate textual modality for guidance. Therefore, MoE retrieval may not be the optimal approach for visual prompt retrieval in LVLMs. It is important to adopt a more robust framework that integrates textual guidance to enhance overall performance effectively.
@@
Retrieval Visualization. We visualize the prompt distribution of retrieval strategies in practical tasks. Taking the MMVet task as an example, Figure 4 presents a bar graph comparing the retrieval decisions of GateNet and our AutoV. The bar heights indicate the frequency of visual prompt retrieval, while the overlaid line plots depict the independent accuracy for each prompt. Our analysis reveals a notable discrepancy between the GateNet retrieval strategy and the actual inference accuracy of visual prompts. For instance, GateNet retrieves , the second-most useful prompt, only times (28/218, of cases), despite its high accuracy. In contrast, AutoV exhibits stronger alignment with the standalone accuracy of visual prompts, and prefers to select and , the two prompts that own superior accuracy.
4.3.2 More candidates perform better?
| Number | Source | MMMU | VizWiz | MMVet | Avg. |
| 0 | Vanilla image | 36.3 | 50.0 | 30.6 | 39.0 |
| 2 | + {L, L} | 37.3 | 50.3 | 32.0 | 39.9 |
| 3 | + {L} | 37.8 | 50.8 | 32.3 | 40.3 |
| 4 | + {L} | 38.3 | 51.0 | 32.8 | 40.7 |
| 5 | + {L} | 38.2 | 50.8 | 32.8 | 40.6 |
| 6 | + {L} | 38.5 | 50.7 | 32.6 | 40.6 |
| 7 | + {RedCircle} | 38.7 | 50.7 | 32.9 | 40.8 |
| 8 | + {FGVP} | 38.7 | 50.4 | 32.4 | 40.5 |
@@ 4.3.2 More candidates perform better?
In AutoV, having more visual prompt candidates can offer a more diverse selection, enhancing the flexibility and adaptability of LVLMs. However, is it always better to have more visual prompts? In the main experiment, we used the attention prompting images [55] from layers , , , and in the CLIP [40] model as inputs, with a total of visual prompts. Here, we consider extra scenarios ranging from prompts and conduct experiments on the MMMU, VizWiz, and MMVet benchmarks. The results are listed in Table 3. First, we observe that when the visual prompts themselves can lead to performance improvement, the accuracy increases as their quantity grows, but the rate of accuracy improvement slows down as the number continues to rise. Taking the experiments on MMMU as an example, as the number of prompts increases, the performance gain remains at when increasing from to prompts. However, when further increasing to prompts, the improvement drops to , even though FGVP alone can contribute a gain. Then, even if a single visual prompt itself degrades accuracy, incorporating it into the candidate pool only marginally harms the overall performance, thanks to the inherent retrieval capability of AutoV. Taking the results on VizWiz as an example: although both RedCircle and FGVP individually show negative gains ( and respectively), including them results in either no overall accuracy drop or just a minimal decrease. In summary, both the quality and quantity of visual prompts impact overall performance. We recommend selecting known high-quality visual prompts as candidates, while limiting to fewer than instances to strike the balance between accuracy and efficiency.
@@ 4.3.3 The transferability of AutoV
The retrieval strategy in our AutoV, trained on one LVLM, can be transferred to various other LVLMs. First, our method demonstrates broad applicability across open-source models. In Table 1, both InternVL2 and Qwen2.5-VL achieve a improvement by directly adopting the retrieval strategy from AutoV pre-trained on LLaVA-OneVision. More remarkably, Table 4 reveals that even closed-source models like Gemini-Pro [46] and GPT-4o [20] show substantial gains, averaging and points respectively. Therefore, AutoV owns good transferability, as its retrieval strategy consistently boosts performance across both open- and closed-source LVLMs without model-specific adaptation.

### v2

@@ 4.2 Main Results
| Method | MMMU | POPE | VizWiz | MMVet | VQA | LLaVA | MMStar | MathVista | Avg. |
| LLaVA-v1.5 (7B) | |||||||||
| Base Model | 36.3 | 85.9 | 50.0 | 30.6 | 58.2 | 65.4 | 33.2 | 34.2 | |
| FGVP (NIPS23) | 36.9 | 72.6 | 47.5 | 29.0 | 46.3 | 55.6 | 33.0 | 34.6 | |
| Circle (ICCV23) | 37.0 | 84.5 | 49.6 | 31.3 | 55.5 | 62.2 | 33.4 | 35.0 | |
| API (ECCV24) | 37.4 | 86.6 | 51.3 | 31.8 | 58.4 | 67.2 | 34.0 | 35.7 | |
| AutoV | 38.7 | 87.4 | 52.1 | 32.9 | 60.6 | 68.6 | 35.2 | 36.9 | |
| LLaVA-v1.5 (13B) | |||||||||
| Base Model | 36.7 | 85.9 | 53.6 | 36.1 | 61.2 | 66.6 | 32.8 | 35.0 | |
| FGVP (NIPS23) | 36.9 | 76.2 | 46.8 | 29.7 | 49.2 | 63.7 | 31.9 | 35.3 | |
| Circle (ICCV23) | 36.9 | 84.6 | 53.8 | 31.2 | 58.1 | 66.9 | 33.5 | 36.2 | |
| API (ECCV24) | 37.3 | 86.8 | 54.6 | 37.6 | 62.1 | 67.6 | 33.0 | 37.0 | |
| AutoV | 38.9 | 87.8 | 56.3 | 38.0 | 62.7 | 69.5 | 34.2 | 38.3 | |
| LLaVA-OneVision (7B) | |||||||||
| Base Model | 49.4 | 89.1 | 58.2 | 48.8 | 76.1 | 80.6 | 61.8 | 63.2 | |
| FGVP (NIPS23) | 49.6 | 82.9 | 58.9 | 49.1 | 65.4 | 70.8 | 58.8 | 63.4 | |
| Circle (ICCV23) | 50.4 | 88.2 | 60.2 | 47.2 | 77.2 | 80.8 | 62.7 | 65.4 | |
| API (ECCV24) | 50.8 | 90.3 | 66.9 | 51.5 | 77.7 | 81.9 | 63.0 | 64.9 | |
| AutoV | 54.0 | 91.0 | 68.4 | 52.9 | 78.0 | 85.3 | 64.5 | 66.8 | |
| InternVL2 (8B) | |||||||||
| Base Model | 48.6 | 87.8 | 58.6 | 46.5 | 77.0 | 74.6 | 62.0 | 58.3 | |
| FGVP (NIPS23) | 49.6 | 84.2 | 58.3 | 44.3 | 70.7 | 75.7 | 60.0 | 58.9 | |
| Circle (ICCV23) | 48.3 | 86.0 | 57.5 | 45.8 | 74.9 | 75.4 | 63.4 | 58.4 | |
| API (ECCV24) | 50.5 | 88.9 | 61.2 | 48.6 | 77.0 | 76.2 | 63.9 | 59.4 | |
| AutoV | 51.7 | 89.6 | 64.8 | 50.2 | 78.0 | 79.5 | 65.8 | 70.5 | |
| Qwen2.5-VL (7B) | |||||||||
| Base Model | 50.1 | 87.5 | 69.5 | 56.6 | 83.0 | 98.0 | 62.4 | 68.2 | |
| FGVP (NIPS23) | 49.6 | 81.6 | 70.8 | 43.1 | 78.0 | 98.2 | 63.4 | 68.0 | |
| Circle (ICCV23) | 51.6 | 84.2 | 69.2 | 44.5 | 80.5 | 97.5 | 65.1 | 69.1 | |
| API (ECCV24) | 51.0 | 87.9 | 72.1 | 58.0 | 85.3 | 100.6 | 63.0 | 67.6 | |
| AutoV | 53.9 | 89.5 | 74.4 | 59.1 | 86.1 | 102.3 | 65.6 | 70.2 | |
Our results are listed in Table 1, and we compare AutoV with three previous visual prompting methods designed for LVLMs, including FGVP [yang2023fine], RedCircle [shtedritski2023does], and API [api]. The training data and implementation details can be found in the Appendix A and C, respectively. Here, we employ the above three types of visual prompts111The pool includes four attention-based visual prompts extracted from different transformer layers of CLIP in API [api], together with two additional prompts of distinct types, resulting in six images in total. as the candidate pool for AutoV to retrieve from optimally.
@@
The four main observations from the experimental results are as follows: (1) Our AutoV significantly outperforms existing methods, demonstrating consistent advantages across all models and benchmarks. Specifically, for LLaVA-7B, LLaVA-13B, LLaVA-OneVision, InternVL2, and Qwen2.5-VL, the average improvements relative to the base model are , , , and , respectively. Our method performs particularly well for LLaVA-OneVision on VizWiz, with an improvement of . Besides, our method appears to be more effective and robust on the VQA task, where other methods often yield negative gains. (2) Our approach shows the most significant performance advantage on MMMU compared to the API. With an average score of for our method versus for theirs, the relative difference of underscores the importance of accurate case-level visual cues retrieval instead of benchmark-level visual prompts selection. The other two visual prompting methods, lacking the ability to adapt to different queries, do not perform well on reasoning tasks. (3) Our method further supplements inherently less accurate models. For instance, LLaVA-7B equipped with our approach achieves higher accuracy than the 13B-scale model on 5 out of 8 tasks. This means that AutoV can be utilized for model efficiency on certain specific tasks: a few additional MLP layers are a better choice compared with hundreds of times larger parameter sizes. (4) Both Qwen2.5-VL and InternVL2 integrate the retrieval optimal prompting strategy from AutoV, which is pre-trained on LLaVA-OneVision. This integration results in an average accuracy gain of . Notably, the improvements are consistent and observed in all individual tasks, including complex reasoning (e.g., MathVista) and vision-centric scenarios (e.g., MMStar). This demonstrates that the effectiveness of the retrieval-enhanced prompting strategy is model-agnostic and generalizes well across heterogeneous LVLM architectures, regardless of pretraining scales or backbone variations.
Additionally, we extend the evaluation to a broader range of vision-language benchmarks, including CVBench, BLINK, Real-World QA, and MMVP, as well as classic tasks such as image captioning, visual grounding, and multimodal classification. We conduct experiments using both LLaVA-OneVision and Qwen2.5-VL as base models to examine the generalization capability of AutoV. The detailed quantitative results are provided in the Appendix D.
4.3 Analysis
4.3.1 Comparison with Other Retrieval Methods
@@
| only | 36.7 | 50.7 | 31.8 | 39.8 |
| only | 36.9 | 51.3 | 31.4 | 39.9 |
| only | 37.4 | 50.6 | 31.7 | 39.9 |
| only | 37.2 | 51.2 | 31.2 | 39.9 |
| Random | 37.1 | 50.7 | 31.3 | 39.7 |
| MoE | 37.6 | 50.7 | 32.0 | 40.1 |
| AutoV | 38.7 | 52.1 | 32.9 | 41.3 |
@@
To analyze why ranking is more suitable for retrieving visual prompts, we compare it with three other retrieval methods. The experiments are conducted on LLaVA-7B. The retrieval pool is identical to that used in the main experiments.
@@
Heuristic Fixed Retrieval. It involves a single fixed type of visual prompts, based on its predefined rules or heuristics. As shown in rows 3 to 6 of Table 2, we find that each visual prompt could potentially be the optimal solution for different benchmarks. For instance, achieves the best results on MMMU, while is optimal on MMVet, and there is a significant variation in the performance of different types of visual prompts across various benchmarks. In summary, heuristic retrieval offers simplicity in retrieval processes with the need for conducting benchmark-level multiple rounds of trials, consuming more resources.
@@
Random Retrieval. It picks up one visual prompt from the candidate pool without any specific criteria or order. The th row of Table 2 lists its results. By comparing with the individual visual prompt, the performance of random retrieval is on par with or slightly superior across the three benchmarks. For instance, the random algorithm outperforms in MMMU by , in MMVet by . It demonstrates that by flexibly selecting from multiple visual prompt candidates, even random retrieval has the potential to outperform individual visual prompts, and the conclusion provides valuable insights for our approach.
@@
MoE (or GateNet) Retrieval. It is a common technique for selecting the most relevant expert during model decision-making, where Gumbel Softmax enables differentiable weighting of the gating network [fedus2022switch]. In our scenario, the features of multiple visual prompts are input to GateNet to determine which one should be retained and fed into the LLM. We insert GateNet into the output of the projection layer and train it with the same amount of data and training time as AutoV. The results can be seen in the th row of Table 2. We observe a improvement with MoE compared to the baseline. However, the improvement compared to the random algorithm was only , lagging significantly behind AutoV by . We suspect that the Mixture-of-Experts architecture encounters a lack of generalization ability during the fine-tuning stage, which can easily lead to overfitting. Therefore, MoE retrieval may not be the optimal approach for visual prompt retrieval in LVLMs. It is important to adopt a more robust framework that integrates textual guidance to enhance overall performance effectively.
@@
Retrieval Visualization. We visualize the prompt distribution of retrieval strategies within - for significance. Figure 4 presents a bar graph comparing the retrieval decisions of GateNet and our AutoV on MMVet. The bar heights indicate the frequency of visual prompt retrieval, while the overlaid line plots depict the independent accuracy for each prompt. Our analysis reveals a notable discrepancy between the GateNet retrieval and the actual inference accuracy of visual prompts. For instance, GateNet retrieves , the second-most useful prompt, only times (38/218, of cases), despite its high accuracy. In contrast, AutoV exhibits stronger alignment with the standalone accuracy of visual prompts, and prefers to select and , the two prompts that own superior accuracy.
4.3.2 Scaling Law in the Visual Prompts Pool
| Pool Size | 0 | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |
| Source | Baseline | +{API-L} | +{API-L} | +{API-L} | +{API-L} | +{RedCircle} | +{FGVP} | +{VTPrompt} | +{TVP} |
| MMMU | 36.3 | 37.3 | 37.5 | 38.3 | 38.3 | 38.6 | 38.7 | 38.7 | 38.9 |
| VizWiz | 50.0 | 50.3 | 50.8 | 51.0 | 51.9 | 52.1 | 52.1 | 52.6 | 52.6 |
| MMVet | 30.6 | 31.8 | 32.0 | 32.6 | 32.8 | 32.9 | 32.9 | 33.1 | 33.1 |
| Avg. | 39.0 | 39.8 | 40.1 | 40.7 | 41.0 | 41.2 | 41.3 | 41.5 | 41.6 |
@@ 4.3.2 Scaling Law in the Visual Prompts Pool
Having more visual prompt candidates can offer a more diverse selection, enhancing the flexibility and adaptability of LVLMs. However, is it always better to have more visual prompts in AutoV? In the main experiment, we used the attention prompting images [api] from layers , , , and in the CLIP [CLIP] model, FGVP [yang2023fine], and RedCircle [shtedritski2023does] as inputs, with a total of visual prompts. Here, we consider extra scenarios ranging from prompts and conduct experiments on the MMMU, VizWiz, and MMVet. The results are listed in Table 3. First, we observe that when the visual prompts themselves can lead to performance improvement, the accuracy increases as their quantity grows, but the rate of accuracy improvement slows down as the number continues to rise. Taking the experiments on MMMU as an example, as the number of prompts increases, the performance gain remains at when increasing from to prompts. However, when further increasing to prompts, the improvement drops to . Then, even if a single visual prompt itself degrades accuracy, incorporating it into the candidate pool does not hurt overall performance and can even lead to slight improvements, thanks to the inherent retrieval capability of AutoV. Taking the results on VizWiz as an example: although both RedCircle and FGVP individually show negative gains ( and , respectively), including them in the candidate pool leads to an overall improvement. In summary, AutoV exhibits notable robustness to variations in visual prompt quality, consistently preserving or enhancing overall performance across diverse settings. We recommend selecting known high-quality visual prompts as candidates, while limiting to fewer than instances to strike the balance between performance and efficiency.
@@ 4.3.3 The Transferability of AutoV
The retrieval strategy in our AutoV, trained on one LVLM, can be transferred to other LVLMs. First, our method demonstrates broad applicability across open-source models. In Table 1, both InternVL2 and Qwen2.5-VL achieve a improvement by directly adopting the retrieval strategy from AutoV pre-trained on LLaVA-OneVision. Besides, Table 4 reveals that even closed-source models like Gemini-Pro [team2024gemini] and GPT-4o [hurst2024gpt] show substantial gains, averaging and points respectively. Therefore, AutoV owns good transferability, as its retrieval strategy consistently boosts performance across both open- and closed-source LVLMs without model-specific adaptation.
4.3.4 Statistical Significance Analysis
To rigorously evaluate the reliability of our reported improvements, we conduct significance testing on the results in Table 1 (LLaVA-1.5 7B) across multiple random seeds. Taking the MMMU benchmark as an example, a paired t-test [hsu2014paired] confirms that improvement of AutoV over the baseline is highly significant (mean , ), with an extremely small -value, indicating that the observed gain is unlikely to arise from random chance. This supports our claim that AutoV consistently enhances VLM performance across repeated runs. Detailed statistical tables and analyses are provided in the Appendix E.